# Antifragile Perimeter Control: Anticipating and Gaining from Disruptions with Reinforcement Learning

### v3

@@ Abstract
The optimal operation of transportation systems is often susceptible to unexpected disruptions, such as traffic accidents and social events. Many established control strategies relying on mathematical models can struggle with real-world disruptions, leading to a significant deviation from their anticipated efficiency. This work applies the cutting-edge concept of antifragility to design a traffic control strategy for urban road networks against disruptions. Antifragility sets itself apart from robustness and resilience as it represents a system’s ability to not only withstand stressors, shocks, and volatility but also to thrive and enhance performance in the presence of such adversarial events. Incorporating antifragile terms composed of traffic state derivatives and redundancy, a model-free deep reinforcement learning algorithm is developed and subsequently evaluated in a two-region cordon-shaped urban traffic perimeter network. Promising results demonstrate (a) the superior performance of the proposed algorithm under both static and incremental magnitude of disruptions compared to the state-of-the-art algorithms, and (b) under incremental disruptions, the proposed method exhibits less fragility compared to the baseline methods. The proposed antifragile methodology has the potential to be integrated into systems subject to disruptions across various disciplines.
@@
antifragility , reinforcement learning , perimeter control , disruptions , macroscopic fundamental diagram
@@ 1 Introduction
Transportation networks serve as vital channels for the movement of people and the flow of goods, and the optimization and control of intelligent transportation systems have led to a multitude of research endeavors and practical implementations (Auer et al., 2016). Given that various sorts of disruptions, such as traffic accidents, social events, and adversarial weather conditions, often occur unexpectedly in real-world networks, examining the robustness and resilience of transportation systems is highly crucial (Ganin et al., 2019). What makes it even more challenging is the continuous growth of motorized traffic. For instance, an approximate 50% increase in traffic demand can be observed in both the U.S. (U.S. Department of Transportation, 2019) and Switzerland (Federal Statistical Office, 2020) over the past few decades. Such growth can lead to not only an escalation of congestion but also more traffic accidents (Dickerson et al., 2000). To picture the consequences, recent work in Sun et al. (2024) has demonstrated the fragile nature of road transportation networks with mathematical proof, demonstrating that performance degrades exponentially with a linearly increasing magnitude of disruptions. Therefore, transportation systems must be designed to secure a decent level of service even when faced with unexpected disruptions of unforeseen magnitudes and exponentially growing negative effects.
@@
To address such issues, the concept of antifragility has shed light on a potential solution. First introduced in the reputable general publication Antifragile: Things That Gain from Disorder (Taleb, 2012), and later mathematically formulated as academic works in Taleb (2013); Taleb and Douady (2013), it has provided insights into designing systems that can benefit from disruptions and perform better under growing volatility and randomness. Its counterpart concept, fragility, can be dated earlier in complexity theory in Vespignani (2010), indicating a cascading effect of interdependent variables in complex networks. Ever since the proposal of antifragility, it has gained substantial interest from both the public and academia, particularly in the complex systems (Axenie et al., 2024) and risk engineering communities (Aven, 2015; Grassi et al., 2024). However, the design principles for realizing antifragile transportation systems are generally unexplored. One promising approach to induce antifragility is applying learning-based algorithms, such as Reinforcement Learning (RL) (Haydari and Yılmaz, 2022), since RL agents can gradually adjust their decision-making when deployed to an environment subject to variations.
@@
The main goal of this paper is to design an antifragile perimeter control algorithm against different types of traffic disruptions with various magnitudes. The key contributions are: a) We formulate and distinguish the concept of antifragility as opposed to other related and commonly used terms in the transportation domain; b) We introduce how antifragility can be incorporated into RL algorithms to achieve superior performance compared to benchmark methods; and c) We propose a skewness-based quantitative indicator to showcase the antifragile properties of our proposed algorithm under environment with increasing disruptions.
@@
The remainder of this paper is structured as follows. In Section 2, we introduce relevant literature on multiple aspects covered in this work, while Section 3 mathematically formulates the cordon-shaped perimeter control simulation environment. Methodologies related to incorporating antifragility into RL algorithms are detailed in Section 4. Section 5 discusses the simulation setups and parametrization. Results are presented in Section 6, followed by conclusion remarks and further discussions in 7.
@@ 2 Literature review
This section reviews the relevant literature with three topics intersecting in this work. First, a macroscopic traffic model and the control strategy applied in this paper are introduced, which serve as the basis of model dynamics for the simulation environment. Next, we study how robustness or resilience can be induced with RL algorithms to counter real-world model uncertainty. Finally, the definitions of robustness, resilience, and antifragile are distinguished while introducing the design philosophy of an antifragile system.
@@ 2.1 Macroscopic fundamental diagram and perimeter control
Perimeter control is among the strategies that have attracted immense attention and research efforts. By mitigating incoming flows from adjacent regions into a protected zone, the traffic density in the protected area remains below the critical density to uphold a satisfactory serviceability. Geroliminis et al. (2013) proposed an optimal perimeter control method using Model Predictive Control (MPC) and proved its effectiveness compared to a greedy controller in a cordon network. Later, Zhou and Gayah (2021) introduced an RL-based algorithm that exhibits superior performance and offers the advantage of being model-free. One major issue with the previous works is the assumption of MFD homogeneity. To tackle this challenge, a substantial amount of effort has been made in investigating the partitioning algorithms so that a well-defined MFD can be created for a sub-network (Ambühl et al., 2019; Saedi et al., 2020), creating a multi-reservoir system for applying perimeter control (Aboudolas and Geroliminis, 2013). However, MFDs in the real world can hardly be well-defined, as demonstrated in Ambühl et al. (2021) with loop detector data over a year. Wang et al. (2015) and Ji et al. (2015) also showed that adverse weather conditions and traffic incidents can alter the shape of the MFDs, and even a recovery from the peak-hour congestion may lead to a hysteresis (Gayah and Daganzo, 2011). These phenomena could potentially violate the mathematical model that serves as the foundation for the established model-based perimeter controllers.
@@ 2.2 Leveraging RL to induce robustness or resilience
Given the presence of real-world disturbances and disruptions, it is essential to assess the robustness and resilience of any newly developed traffic control algorithm. As RL offers flexibility in defining state representations, actions, and rewards, researchers have leveraged these features to enhance the robustness and resilience of RL-based traffic control strategies. Here, we present a collection of studies that apply RL in traffic control and claim to demonstrate robustness or resilience. While Zhou and Gayah (2021); Chen et al. (2022); Su et al. (2023); Zhou and Gayah (2023) focus on perimeter control frameworks, Aslani et al. (2018); Rodrigues and Azevedo (2019); Chu et al. (2020); Tan et al. (2020); Wu et al. (2020); Korecki et al. (2023) explore the design for signalized intersections. Some of these works directly compare the performance of the proposed algorithm with the benchmark methods under disruptions to showcase either robustness or resilience, whereas Rodrigues and Azevedo (2019); Tan et al. (2020); Chu et al. (2020); Zhou and Gayah (2023) further state explicitly how it is induced with modifications of the algorithm. For instance, Rodrigues and Azevedo (2019) induced robustness by adding the elapsed time since the last green signal for each phase, Tan et al. (2020) experimented with speed or residual queue as an additional state representation in the RL algorithm, Chu et al. (2020) supplemented the control policies of neighboring intersections as additional information to the agents, and Zhou and Gayah (2023) used an extra binary congestion indicator in the state space. As a result, the algorithms are given additional information related to disturbances or disruptions of the environment. The analysis of the state-of-the-art RL studies considers reversals or sudden changes in the state-action-reward dynamics, which evokes unanticipated uncertainty. The problem in these contexts is often to respond to unexpected results appropriately since they might indicate a shift in the environment. In this case, exploration refers to the process of looking for new information to improve the RL agent’s understanding of the traffic dynamics under disruptions, which would then be used to identify better courses of action. By exploring how robustness and resilience can be achieved using RL-based traffic control algorithms, we can potentially develop antifragile traffic control systems through similar approaches.
@@ 2.3 Antifragility: definitions and distinctions
Before diving into antifragile system design, we first need to review the distinctions between robustness, resilience, and antifragility. We follow the definitions proposed by Zhou et al. (2019), that robustness is concerned with assessing a system’s capacity to preserve its initial state and resist performance deterioration under minor disturbances, while resilience emphasizes the ability and speed of a system to recover from major disruptions to the original state. In Taleb (2012), two levels of antifragility have been indicated: proto-antifragility and antifragility. To differentiate these two terms, we rename the latter and the more important one as progressive antifragility. Proto-antifragility describes systems that can improve performance when exposed to moderate disruptions with magnitudes in a similar range, resembling the concept of adaptiveness. An illustrated example can be the biological process of hormesis. The higher level of antifragility, i.e., progressive antifragility, emphasizes the concave response of the system under an increasing magnitude of disruptions, which can be mathematically formulated with Jensen’s inequality $\mathbb{E}[g(X)]\leq g(\mathbb{E}[X])$. A graphical comparison between robustness, resilience, adaptiveness, and antifragility is shown in Fig. 13. Although Sun et al. (2024) has proven mathematically that urban road networks are progressively fragile by nature, it should be noted that they can be less fragile when proper control strategies are adopted to induce antifragility.
@@
Ever since the concept of antifragility was proposed, it has become an increasingly popular concept across various disciplines, such as energy (Coppitters and Contino, 2023), biology (Kim et al., 2020), medicine (Axenie et al., 2022), and robotics (Axenie and Saveriano, 2023). Researchers have also proposed methods to incentivize the antifragile property of a system by emphasizing the derivatives to capture the temporal evolution patterns of the system dynamics, i.e., how fast the system state deviates toward possible black swan events and the curvature of this deviation (Taleb and Douady, 2013; Taleb and West, 2023; Axenie et al., 2022). With this additional information, the system can anticipate ongoing disruptions and be more responsive to drastic changes. Similar to the function of redundancy in resilience (Tan et al., 2019; Kamalahmadi et al., 2022), redundancy can also be added to the system to induce antifragility de Bruijn et al. (2020); Johnson and Gheorghe (2013); Munoz et al. (2022). Other feasible approaches also include time-scale separation and attractor dynamics (Axenie et al., 2022). However, current studies on antifragility in engineering mostly pertain to proto-antifragility, such as post-disaster treatment and reconstruction efforts, as in Fang and Sansavini (2017); Priyadarshini et al. (2022). Leveraging the potential of antifragility for the operation of transportation systems is still a largely novel and unexplored notion.
@@ 3 Problem Formulation
This paper studies perimeter control between two homogeneous cordon-shaped urban networks, as in Geroliminis et al. (2013), with the inner region representing a city center, as shown in Fig. 1(a). Additional traffic demand at time $t$ for an Origin-Destination (OD) pair from region $i$ to region $j$ is denoted as $q_{ij}(t)$. The inner and outer regions are assumed to have different MFDs to represent the capacity difference to accommodate vehicles within the network between the city center and the surrounding region, which are defined as $G_{i}(n_{i}(t))$ as illustrated in Fig. 1(b). Given the total number of vehicles with presence in region $i$ at time $t$, denoted as $n_{i}(t)$, the total trip completion rate for this region $i$, denoted as $M_{i}(t)$, can be determined using the corresponding MFD, which comprises both the intraregional trip completion, i.e., $M_{ii}(t)$ and the interregional transfer flow, i.e., $M_{ij}(t)\>(i\neq j)$ with $i,j\in\{1,2\}$. To protect regions from being overflown by high traffic demand, the percentage of the transfer flow $M_{ij}(t)$ allowed to go across the perimeter at time $t$ is regulated by two perimeter controllers denoted as $u_{ij}(t)\>(i\neq j)$. A list of all notations used in this paper, including the notations used in defining the RL algorithm and the antifragile terms, is summarized in Table 1.
@@ 3. Notations for the additional antifragile terms applied in reinforcement learning
Eq. 1a describes the change rate of the intraregional vehicle accumulation of region $i$. It is the sum of intraregional traffic demand in this region, denoted as $q_{ii}(t)$, together with the perimeter control regulated transfer flow from region $j$ to region $i$, defined as $u_{ji}(t)\cdot M_{ji}(t)$, then deducted by the trip completion within region $i$, denoted as $M_{ii}(t)$. similarly, the change rate of interregional traffic accumulation, as in Eq. 1b shows, is the difference between the interregional traffic demand, denoted as $q_{ij}(t)$, and the regulated transfer flow $u_{ij}(t)\cdot M_{ij}(t)$:
@@
The total trip completion, i.e., $M_{i}(t)$ for region $i$ at time $t$ is calculated based on the trip accumulation and the related MFD, defined as $G_{i}(n_{i}(t))$, and is the sum of the intraregional trip completion, i.e., $M_{ii}(t)$, in Eq. 2a and the interregional transfer flow, i.e., $M_{ij}(t)\>(i\neq j)$, in Eq. 2b:
@@
The objective function is to maximize the throughput of this cordon-shaped network, which is the sum of the intraregional trip completion in both regions.
@@
Intraregional and interregional vehicle accumulation $n_{ii}(t)$ and $n_{ij}(t)$ are non-negative values, and $n_{i,\rm cap}$ is the maximal possible number of vehicles accumulated in the region $i$, at which value a gridlock will occur in the network. $u_{\min}$ and $u_{\max}$ represent the lower and upper limits for the perimeter control variable $u_{ij}(t)$ for both directions. Such boundaries are in line with Geroliminis et al. (2013); Zhou and Gayah (2021), as perimeter control is normally implemented through signalization. While $u_{\max}$ accounts for the lost time between the red and green phases, $u_{\min}$ is necessitated since an indefinite long red light is rare in real-world urban networks.
In contrast to the objective function to be applied for control-based strategies, the objective function for the proposed antifragile RL-based algorithms $J_{anti}$ is:

### v4

@@ Abstract
The optimal operation of transportation systems is often susceptible to unexpected disruptions, such as traffic accidents and social events. Many established control strategies relying on mathematical models can struggle with real-world disruptions, leading to a significant deviation from their anticipated efficiency. This work applies the cutting-edge concept of antifragility to design a traffic control strategy for urban road networks against disruptions. Antifragility sets itself apart from robustness, resilience, and reliability as it represents a system’s ability to not only withstand stressors, shocks, and volatility but also to thrive and enhance performance in the presence of such adversarial events. Incorporating antifragile terms composed of traffic state derivatives and redundancy, a model-free deep reinforcement learning algorithm is developed and subsequently evaluated in a two-region cordon-shaped urban traffic perimeter network. Promising results highlight (a) the superior performance of the proposed algorithm compared to the state-of-the-art methods under incremental magnitude of disruptions, (b) distribution skewness as the antifragility indicator demonstrating its relative antifragility, and (c) its effectiveness under limited observability due to real-world data availability constraints. The proposed antifragile methodology is generalizable and holds potential for application beyond perimeter control, offering integration into systems exposed to disruptions across various disciplines.
@@
antifragility , reinforcement learning , perimeter control , demand and supply disruption , macroscopic fundamental diagram
@@ 1 Introduction
Transportation networks serve as vital channels for the movement of people and the flow of goods, and the optimization and control of intelligent transportation systems have led to a multitude of research endeavors and practical implementations (Auer et al., 2016). Given that various sorts of disruptions, such as traffic accidents, social events, and adversarial weather conditions, often occur unexpectedly in real-world networks, examining the robustness and resilience of transportation systems is highly crucial (Ganin et al., 2019). What makes it even more challenging is the continuous growth of motorized traffic. For instance, an approximate 50% increase in traffic demand can be observed in both the U.S. (U.S. Department of Transportation, 2019) and Switzerland (Federal Statistical Office, 2020) over the past few decades. Such growth can lead to not only an escalation of congestion but also more traffic accidents (Dickerson et al., 2000). To picture the consequences, recent work in Sun et al. (2024) has demonstrated the fragile nature of road transportation networks with mathematical proof, demonstrating that performance degrades exponentially with a linearly increasing magnitude of disruptions. Therefore, transportation systems must be designed to secure a decent level of service even when faced with unexpected disruptions of unforeseen magnitudes and the exponentially growing negative effects.
@@
To address such issues, the concept of antifragility has shed light on a potential solution. First introduced in the reputable general publication Antifragile: Things That Gain from Disorder (Taleb, 2012), and later mathematically formulated as academic works in Taleb (2013); Taleb and Douady (2013), antifragility has provided insights into designing systems that can benefit from disruptions and perform better under growing volatility and randomness. Its counterpart concept, fragility, can be dated earlier in complexity theory in Vespignani (2010), indicating a cascading effect of interdependent variables in complex networks, such as in transportation networks (Cats and Hijner, 2021). Ever since the concept was proposed, antifragility has gained substantial interest from both the public and academia, particularly in the complex systems (Axenie et al., 2024) and risk engineering communities (Aven, 2015; Thekdi and Aven, 2019; Grassi et al., 2024). However, the design principles for realizing antifragile transportation systems are generally unexplored. One promising approach to induce antifragility is applying learning-based algorithms, such as Reinforcement Learning (RL) (Haydari and Yılmaz, 2022), since RL agents can gradually adjust their decision-making when deployed to an environment subject to variations.
@@
The main goal of this paper is to design an antifragile perimeter control algorithm capable of adapting to increasing magnitudes of disruption, reflecting the effects of urban densification and rising traffic demand. The key contributions are: a) We formulate and distinguish the concept of antifragility as opposed to other related and commonly used terms in the transportation domain; b) We introduce how antifragility can be incorporated into RL algorithms to achieve superior performance compared to benchmark methods, tackling both fragile performance and observability issues; and c) We adopt a skewness-based quantitative indicator to showcase the antifragile properties of our proposed algorithm under increasing disruptions. A contribution map following previous works is illustrated in Fig. 1.
Figure 1: Map of contribution
@@
The remainder of this paper is structured as follows. In Section 2, we introduce relevant literature on multiple aspects covered in this work, while Section 3 mathematically formulates the cordon-shaped perimeter control simulation environment. Methodologies related to incorporating antifragility into RL algorithms are detailed in Section 4. Section 5 discusses the simulation setups and parametrization. Results are presented in Section 6, followed by concluding remarks and further discussions in 7.
@@ 2 Literature review
This section reviews the relevant literature on three topics intertwined in this work. First, a macroscopic traffic model and the control strategy applied in this paper are introduced, which serve as the basis of model dynamics for the simulation environment. Next, we study how robustness or resilience can be induced with RL algorithms to counter real-world model uncertainty. Finally, related terminologies to antifragility are distinguished while introducing the design philosophy of an antifragile system.
@@ 2.1 Macroscopic fundamental diagram and perimeter control
Perimeter control is among the strategies that have attracted immense attention and research efforts. By mitigating incoming flows from adjacent regions into a protected zone, the traffic density in the protected area remains below the critical density to uphold a satisfactory serviceability. Geroliminis et al. (2013) proposed an optimal perimeter control method using Model Predictive Control (MPC) and proved its effectiveness compared to a greedy controller in a cordon network. Later, Sirmatel and Geroliminis (2020) introduced a Moving Horizon Estimation (MHE) scheme together with MPC to tackle further measurement noise and the observability issues. One major issue of the previous works is the assumption of MFD homogeneity, and a substantial amount of effort has been made in investigating the partitioning algorithms so that a well-defined MFD can be created for each sub-network (Ambühl et al., 2019; Saedi et al., 2020), creating a multi-reservoir system for applying perimeter control (Aboudolas and Geroliminis, 2013). However, MFDs in the real world can hardly be well-defined, as demonstrated in Ambühl et al. (2021) with loop detector data over a year. Wang et al. (2015) and Ji et al. (2015) also showed that adverse weather conditions and traffic incidents can alter the shape of the MFDs, and even a recovery from the peak-hour congestion may lead to a hysteresis (Gayah and Daganzo, 2011). These phenomena could potentially violate the mathematical model that serves as the foundation for the established model-based perimeter controllers. Therefore, Zhou and Gayah (2021) introduced an RL-based algorithm that exhibits superior performance and offers the advantage of being model-free.
@@ 2.2 Leveraging RL to induce robustness or resilience
Given the presence of real-world disturbances and disruptions, it is essential to assess the robustness and resilience of any newly developed traffic control algorithm. As RL offers flexibility in defining state representations, actions, and rewards, researchers have leveraged these features to enhance the robustness and resilience of RL-based traffic control strategies. Here, we present a collection of studies that apply RL in traffic control and claim to demonstrate robustness or resilience. While Zhou and Gayah (2021); Chen et al. (2022); Su et al. (2023); Zhou and Gayah (2023) focus on perimeter control frameworks, Aslani et al. (2018); Rodrigues and Azevedo (2019); Chu et al. (2020); Tan et al. (2020); Wu et al. (2020); Korecki et al. (2023) explore the design for signalized intersections. Some of these works directly compare the performance of the proposed algorithm with the benchmark methods under disruptions to showcase either robustness or resilience, whereas Rodrigues and Azevedo (2019); Tan et al. (2020); Chu et al. (2020); Zhou and Gayah (2023) further mention explicitly how it is induced with modifications of the algorithm. For instance, Rodrigues and Azevedo (2019) induced robustness by adding the elapsed time since the last green signal for each phase, Tan et al. (2020) experimented with speed or residual queue as an additional state representation in the RL algorithm, Chu et al. (2020) supplemented the control policies of neighboring intersections as additional information to the agents, and Zhou and Gayah (2023) used an extra binary congestion indicator in the state space. As a result, the algorithms are given additional information related to disturbances or disruptions of the environment. The analysis of the state-of-the-art RL studies considers reversals or sudden changes in the state-action-reward dynamics, which evoke unanticipated uncertainty. The problem in these contexts is often to respond to unexpected results appropriately, since they might indicate a shift in the environment. In this case, exploration refers to the process of looking for new information to improve the RL agent’s understanding of the traffic dynamics under disruptions, which would then be used to identify better courses of action. By exploring how robustness and resilience can be achieved using RL-based traffic control algorithms, we can potentially develop antifragile traffic control systems through similar approaches.
@@ 2.3 Antifragility: definitions and distinctions
Before diving into antifragile system design, we first need to review the distinctions between antifragility and other closely related terms, which include robustness, resilience, reliability, and adaptiveness. We follow the definitions proposed in Zhou et al. (2019), that robustness is concerned with assessing a system’s capacity to preserve its initial state and resist performance deterioration under minor disturbances, while resilience emphasizes the ability and speed of a system to recover from major disruptions to the original state. Reliability in transportation has a much broader range of meanings. Pennetti et al. (2020) proposed travel time reliability, which has a similar flavor to robustness in measuring deviation from normal operation, but focuses on the probability of such deviation exceeding a certain threshold. Cats et al. (2017) introduced a reliability assessment indicator to account for performance loss for the entire range of possible capacity reductions, whereas robustness pertains merely to a single disruption event. Adaptiveness describes the ability of systems to alter their traits to satisfy autonomy in different environments (Hooker, 2011), which aligns with the concept of proto-antifragility proposed in Taleb (2012). However, adaptiveness does not necessarily guarantee a consistent performance improvement under growing magnitudes of disruptions. Therefore, Taleb (2012) introduced the concept of antifragility, emphasizing the concave response of the system under increasing disruptions, which can be mathematically formulated with Jensen’s inequality $\mathbb{E}[g(X)]\leq g(\mathbb{E}[X])$. A graphical comparison between robustness, resilience, adaptiveness, and antifragility is shown in Fig. 16. It should be noted that while (anti-)fragility can be an innate property of a system, proper intervention and control strategies can lead the system to be more antifragile (Axenie et al., 2024). For instance, building on the mathematical proof by Sun et al. (2024) that urban road networks are naturally fragile, this paper seeks to develop antifragile solutions that mitigate such fragility through perimeter control.
@@
Ever since the concept of antifragility was proposed, it has become an increasingly popular concept across various disciplines in recent years, such as energy (Coppitters and Contino, 2023), electricity (Rachunok and Nateghi, 2020), biology (Kim et al., 2020), medicine (Axenie et al., 2022), cyber-systems (Chatterjee and Thekdi, 2020), and robotics (Axenie and Saveriano, 2023). Researchers have also proposed methods to incentivize the antifragile property of a system by emphasizing the derivatives of system states to capture the temporal evolution patterns of the system dynamics, i.e., how fast the system state deviates toward possible black swan events and the curvature of this deviation (Taleb and Douady, 2013; Taleb and West, 2023; Axenie et al., 2022). With this additional information, the system can anticipate ongoing disruptions and be more proactive under drastic changes. Similar to its function in resilience (Tan et al., 2019; Kamalahmadi et al., 2022), redundancy can also be added to the system to induce antifragility de Bruijn et al. (2020); Johnson and Gheorghe (2013); Munoz et al. (2022). Other feasible approaches also include time-scale separation and attractor dynamics (Axenie et al., 2022). However, leveraging the potential of antifragile system design for the operation of transportation networks is still a largely novel and unexplored notion.
@@ 3 Problem formulation
This paper studies perimeter control between two homogeneous cordon-shaped urban networks, as in Geroliminis et al. (2013), with the inner region representing a city center, as shown in Fig. 2(a). The total number of vehicles in region $i$ at time $t$ is denoted as $n_{i}(t)$, while the Origin-Destination (OD) pair can be further divided as $n_{ij}(t)$. The inner and outer regions are assumed to have different MFDs to represent the capacity difference of accommodating vehicles within the network between the city center and the surrounding region, which are defined as $G_{i}(n_{i}(t))$ as illustrated in Fig. 2(b). The total trip completion rate $M_{i}(t)$ for region $i$ at time $t$ can be determined through the corresponding MFD, which comprises both the intraregional trip completion $M_{ii}(t)$ and the interregional transfer flow $M_{ij}(t)\>(i\neq j)$ with $i,j\in\{1,2\}$. While $q_{i}$ denotes the new demand within region $i$ at time $t$, $q_{ij}(t)$ represents the additional demand based on an Origin-Destination (OD) pair from region $i$ to region $j$. To protect regions from being overflown by high traffic demand, the percentage of the transfer flow $M_{ij}(t)$ allowed to go across the perimeter at time $t$ is regulated by two perimeter controllers, denoted as $u_{ij}(t)\>(i\neq j)$ with $i,j\in\{1,2\}$. A list of all notations used in this paper, including the notations used in defining the RL algorithm and the antifragile terms, is summarized in Table 1.
@@ 3. Notations for the additional antifragile terms applied in reinforcement learning
Eq. 1a describes the change rate of the intraregional vehicle accumulation of region $i$. It is the sum of intraregional traffic demand in this region $q_{ii}(t)$ with the perimeter control regulated transfer flow $u_{ji}(t)\cdot M_{ji}(t)$, then deducted by trip completion $M_{ii}(t)$. similarly, the change rate of interregional traffic accumulation, as Eq. 1b shows, is the difference between the interregional traffic demand $q_{ij}(t)$ and the regulated transfer flow $u_{ij}(t)\cdot M_{ij}(t)$:
@@
The total trip completion $M_{i}(t)$ for region $i$ at time $t$ is calculated based on the trip accumulation and the related MFD $G_{i}(n_{i}(t))$, which is the sum of the intraregional trip completion $M_{ii}(t)$ and the interregional transfer flow $M_{ij}(t)\>(i\neq j)$:
@@
The objective function is to maximize the throughput of this cordon-shaped network, which is the sum of the intraregional trip completion $M_{ii}(t)$ in both the inner and the outer regions, while the interregional trip completion represents only vehicles finishing part of their trips.
@@
Intraregional and interregional vehicle accumulation $n_{ii}(t)$ and $n_{ij}(t)$ are non-negative values, and $n_{i,\rm cap}$ is the maximal possible number of vehicles accumulated in region $i$, at which value a gridlock will occur in the network. $u_{\min}$ and $u_{\max}$ represent the lower and upper limits for the perimeter control variable $u_{ij}(t)$ for both directions. Such boundaries are in line with Geroliminis et al. (2013); Zhou and Gayah (2021), as perimeter control is normally implemented through signalization. While $u_{\max}$ accounts for the lost time between the red and green phases, $u_{\min}$ is essential since an indefinitely long red light is rare in real-world urban networks.
In contrast to the objective function to be applied for control-based strategies, the reward function for the proposed antifragile RL-based algorithms $J_{anti}$ is in Eq. 4, with $r_{com}(t)$ standing for the same trip completion as in Eq. 3, i.e., $r_{com}(t)=\sum_{i=1,2}M_{ii}(t)$.